# Quasinormal Modes and Dynamical Evolution of Scalar Fields in the Einstein–Bumblebee Theory with a Cosmological Constant

### v1

@@ Abstract
This paper investigates the dynamic behavior of static, spherically symmetric black holes within the Einstein–Bumblebee gravity model with a cosmological constant, focusing on scalar field perturbations. Through separation of the angular components, the scalar field perturbations outside the black hole are reduced to a purely radial main equation. The quasinormal modes (QNMs) of the system are then determined via the WKB approximation in the frequency domain, while the dynamic evolution of the system is examined in the time domain using finite difference methods. The eigenfrequencies of the waveforms from the time-domain evolution are fitted to cross-validate the frequency-domain results. The study finds that the Lorentz violation parameter $\ell$ and the cosmological constant $\Lambda$ significantly influence the QNMs. Specifically, as $\ell$ increases, the real and imaginary components of the lower modes decrease, while in higher modes, the real part changes minimally, and the imaginary part decreases rapidly. An increase in $\Lambda$ similarly results in a decrease in the overall QNM values. These results are supported by the time-domain analysis, providing a clearer picture of how Lorentz symmetry breaking affects the QNMs of de Sitter spacetime.
Keywords: Lorentz violation; Black hole; Scalar perturbation; Quasinormal modes;
@@ I Introduction
In the context of general relativity, Lorentz invariance is a fundamental principle that underpins both general relativity and the standard model of particle physics Magueijo and Smolin (2002); Li et al. (2011); Heisenberg (2019); Martínez-Huerta et al. (2020). However, at high energy scales, particularly within quantum gravity, Lorentz invariance may no longer hold Ashtekar and Bianchi (2021); Campiglia et al. (2017). This limitation has prompted a growing interest in investigating Lorentz violation as a key path toward deepening our understanding of the fundamental principles of modern physics. The standard model extension is an effective field theory that integrates general relativity with the standard model, incorporating dynamic curvature modes and terms that account for Lorentz-violating effects at the Planck scale Colladay and Kostelecky (1998). Under the influence of the local Lorentz force, violation of Lorentz symmetry can be detected by nonzero vacuum expectation values of these indices Roberts (2017); Yang et al. (2023); Duan et al. (2024); Filho et al. (2024a); Liu et al. (2024a, b); Araújo Filho et al. (2024a); Hosseinifar et al. (2024); Chen et al. (2024). A notable approach to exploring Lorentz violation is the Bumblebee model, a variant of Einstein’s aether theory, characterized by a potential function that minimally rolls towards its vacuum expectation values Berglund et al. (2013). Initially proposed by Kostelecky and Samuel in 1989 as a model for spontaneous Lorentz-symmetry breaking Kostelecky and Samuel (1989), this model has since been studied for its implications in various contexts. Casana et al. identified an exact solution in the Bumblebee gravity model that resembles the Schwarzschild solution, examining its effects on general relativistic phenomena such as planetary perihelion progression, light bending, and time delays caused by curvature Casana et al. (2018). Additionally, the solution has been extended to incorporate a cosmological constant by Maluf Maluf and Neves (2021), a significant advancement that enables a deeper exploration of the effects of Lorentz violation in both de Sitter and anti-de Sitter spacetimes. Given the accelerated expansion of the universe, current models suggest that the universe is asymptotically de Sitter. Thus, this paper aims to explore the effects of Lorentz violation on the dynamical properties of black holes in an asymptotically de Sitter spacetime. Furthermore, within this framework, more black hole solutions have been solved Filho et al. (2023); Övgün et al. (2019); Ding et al. (2020, 2023, 2022). Since then, this solution has served as a foundation for further research into related topics, including quasinormal modes (QNMs) Oliveira et al. (2021); Liu et al. (2023a); Malik (2024); Mai et al. (2024), black hole shadows Jha and Rahaman (2021); Kuang and Övgün (2022); Gao (2024a); Pantig et al. (2024); Liu et al. (2024c, d, e); Afrin et al. (2024), photon trajectories Lambiase et al. (2023a), and studies on entanglement degradation Liu et al. (2025) or particle creation Araújo Filho (2025), as well as the metric affine case Filho et al. (2024b); Amarilo et al. (2024); Lambiase et al. (2023b); Araújo Filho et al. (2024b); Gao (2024b); Heidari et al. (2024); Jha and Rahaman (2024), among others Araújo Filho et al. (2025); Filho et al. (2024c); Liu et al. (2024f); Ji et al. (2024).
@@
On the other hand, in black hole physics, interactions between black holes and matter fields are crucial for understanding black hole stability, testing the no-hair theorem Bhattacharya and Lahiri (2007), and developing new black hole solutions Herdeiro et al. (2016). To study these interactions, researchers often use the linear approximation of black hole perturbation theory Calcagni and Modesto (2017); Mukohyama et al. (2022). Black hole perturbations can be divided into two types. The first type involves perturbations of the black hole’s metric itself, where gravitational radiation from black hole disturbances exceeds that caused by external fields Del Águila and Matos (2021), making it a key measure of black hole stability. The second type involves perturbations within the ambient spacetime of the black hole, arising from disturbances caused by matter fields. These perturbations reflect the dynamic behavior of matter fields near the black hole and offer insights into the interactions between black holes and their environment. Similar to ripples on water caused by external forces, spacetime oscillations propagate as gravitational waves, with energy gradually dissipating during transmission. Perturbative fields in black hole spacetimes typically undergo three stages: initial bursts, QNMs, and power-law tails Destounis et al. (2020). The latter two stages, linked to the intrinsic properties of the black hole, are of particular interest. These stages, often described as the ”ringing” of black holes, exhibit characteristics akin to the acoustic signatures of black holes Chakrabarti and Tchrakian (2002). By analyzing these signatures, researchers can identify and categorize different types of black holes.
@@
This paper investigates the QNMs and the dynamic waveform evolution in de Sitter black hole spacetimes within the Einstein–Bumblebee gravitational model and contrasts these outcomes with the results from traditional Schwarzschild-de Sitter spacetimes Zhidenko (2004). The impact of different Lorentz violation parameters on the properties of black hole spacetimes is explored. Sec. II of this manuscript provides a brief overview of the Einstein–Bumblebee gravity model and its underlying geometrical background and derives equations of motion for the scalar field. Sec. III describes the WKB method Konoplya (2003), employs numerical techniques to find the scalar field QNMs, and presents the associated numerical results. In Sec. IV, we delve into the dynamical evolution of the scalar field under a small cosmological constant, revealing significant late-time behavior characterized by power-law tails. The final section offers a perspective on potential future research directions for this topic.
II Einstein–Bumblebee Gravity Model
@@
Here, $T_{\mu\nu}^{M}$ is the energy–momentum tensor of the matter field and
@@
is the Bumblebee energy–momentum tensor with
@@
By solving this system of differential equations, one can obtain curved spacetime metrics with asymptotically de Sitter (dS) or anti-de Sitter (AdS) properties under the influence of the Bumblebee field. These can also be referred to as Einstein–Bumblebee black hole metrics, and their line element is given by:
@@
The metric function in references can thus be rewritten as $F(r)=1-\frac{2M}{r}-\frac{\Lambda r^{2}}{3}$, which is consistent with the Schwarzschild–de Sitter case. However, the asymptotic behavior of the metric at infinity is not consistent with the Schwarzschild–de Sitter case, and the influence of Lorentz violation also needs to be considered. When the cosmological constant $\Lambda$ is zero, the resulting metric coincides with the results obtained by Casana in 2018 Casana et al. (2018). When the Lorentz violation parameter $\ell=0$, the metric simplifies to Schwarzschild–(A)dS. The incorporation of the cosmological constant modifies the structure of spacetime, resulting in a dual horizon configuration that encompasses both the black hole event horizons and cosmological horizons. This means that, the metric function can be reformulated as Jing (2004); Liu et al. (2023b, 2024g):
@@
II.1 Scalar Field Perturbation Equation
The equation of motion for a massless scalar field is described by the Klein–Gordon equation:
$\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Phi\right)=0,$ (9)
@@ II.1 Scalar Field Perturbation Equation
$V(r)=$ $\frac{F(r)}{r^{2}}\left[(1+\ell)L(L+1)+\frac{2M}{r}-\frac{2\Lambda r^{2}}{3}\right].$ (13)
@@
This equation is a standard wave equation, which means that in our subsequent work, we can solve it via conventional methods, such as the WKB approximation Iyer and Will (1987). Meanwhile, we plotted the potential $V(r)$ against $r$ for various parameters in Fig. 1. From the potential plot, we can see that it satisfies the criteria required by the WKB method.
@@
In such black hole spacetime systems, the dissipation of energy primarily arises from the imposition of fully absorbing boundary conditions. Specifically, this includes two requirements: at the black hole horizon, the perturbation must be an incoming wave, and at infinity, it should be an outgoing wave. These boundary conditions apply to asymptotically flat spacetimes. More precisely, the boundary conditions encompass two stipulations: perturbations must manifest as ingoing waves at the black hole horizon and as outgoing waves at infinity. However, in de Sitter or anti-de Sitter spacetimes, the choice of boundary conditions differs. In de Sitter spacetimes, the de Sitter horizon replaces the role traditionally played by spatial infinity; in anti-de Sitter spacetimes, the requirement for the convergence of solutions at spatial infinity supersedes the condition for outgoing waves. These distinct boundary conditions reflect the unique properties of different types of spacetimes. If equations in the frequency domain are considered, one can separate variables in the wave function, allowing for:
$\psi(t,r)=\psi(r)e^{-i\omega t}.$ (17)
At this point, the boundary conditions for the wave equation are:
$\psi(r)\sim\begin{cases}e^{-i\sqrt{1+\ell}\omega r_{*}},&r\to r_{h},\\ e^{i\sqrt{1+\ell}\omega r_{*}},&r\to r_{c},\end{cases}$ (18)
where $r_{h}$ and $r_{c}$ correspond to the vicinity of the black hole horizon and the de Sitter horizon, respectively. The quasinormal mode problem fundamentally represents a frequency eigenvalue problem for the wave equation. The real and imaginary parts of the quasinormal mode frequency $\omega=\omega_{R}+i\omega_{I}$ each have distinct physical meanings expressed by $e^{-i\omega t}=e^{-i\omega_{R}t}e^{\omega_{I}t}$. The real part $\omega_{R}$ clearly represents the oscillation frequency of the perturbation mode, whereas the imaginary part $\omega_{I}$ indicates the rate of exponential growth or decay of the perturbation. If $\omega_{I}$ is negative, it indicates that the perturbation decays over time, suggesting that the mode is stable; conversely, a positive $\omega_{I}$ signifies an unstable mode, under which the black hole spacetime might undergo substantial changes.
@@ III Numerical Methods and Quasinormal Modes
III.1 Wenzel–Kramers–Brillouin (WKB) Approximation Method

### v2

@@ Abstract
In this paper, we investigate the quasi normal modes (QNMs) of static spherically symmetric black holes in the Einstein—Bumblebee gravity model, taking into account the effects of the field mass and the cosmological constant. Through separation of the angular components, the scalar field perturbations outside the black hole are reduced to a purely radial main equation. We calculated the quasinormal modes using the matrix method and the WKB approximation, and also studied the dynamical evolution of the purely radial main equation using the finite difference method in the time domain. The eigenfrequencies of the waveforms from the time-domain evolution are fitted to cross-validate the frequency-domain results. The Lorentz-violating parameter $\ell$, cosmological constant $\Lambda$, and scalar field mass $\mu M$ affect QNMs. Specifically, increasing $\ell$ decreases both real and imaginary parts of the monopole modes, but in the dipole and quadrupole modes, the real part remains nearly unchanged while the imaginary part drops rapidly; rising $\Lambda$ reduces both parts of QNMs; increasing $\mu M$ raises the real part and lowers the imaginary part. Time-domain analysis confirms these findings, clarifying how Lorentz symmetry breaking impacts QNMs in ds Sitter spacetime.
Keywords: Lorentz violation; Black hole; Quasinormal modes;
@@ I Introduction
In the context of general relativity, Lorentz invariance is a fundamental principle that underpins both general relativity and the standard model of particle physics Magueijo and Smolin (2002); Li et al. (2011); Heisenberg (2019); Martínez-Huerta et al. (2020). However, at high energy scales, particularly within quantum gravity, Lorentz invariance may no longer hold Ashtekar and Bianchi (2021); Campiglia et al. (2017). This limitation has prompted a growing interest in investigating Lorentz violation as a key path toward deepening our understanding of the fundamental principles of modern physics. The standard model extension is an effective field theory that integrates general relativity with the standard model, incorporating dynamic curvature modes and terms that account for Lorentz-violating effects at the Planck scale Colladay and Kostelecky (1998). Under the influence of the local Lorentz force, violation of Lorentz symmetry can be detected by nonzero vacuum expectation values of these indices Roberts (2017); Yang et al. (2023); Duan et al. (2024); Filho et al. (2024a); Liu et al. (2024a, b); Araújo Filho et al. (2024a); Hosseinifar et al. (2024); Chen et al. (2024). A notable approach to exploring Lorentz violation is the Bumblebee model, a variant of Einstein’s aether theory, characterized by a potential function that minimally rolls towards its vacuum expectation values Berglund et al. (2013). Initially proposed by Kostelecky and Samuel in 1989 as a model for spontaneous Lorentz-symmetry breaking Kostelecky and Samuel (1989), this model has since been studied for its implications in various contexts. Casana et al. identified an exact solution in the Bumblebee gravity model that resembles the Schwarzschild solution, examining its effects on general relativistic phenomena such as planetary perihelion progression, light bending, and time delays caused by curvature Casana et al. (2018). Additionally, the solution has been extended to incorporate a cosmological constant by Maluf Maluf and Neves (2021), a significant advancement that enables a deeper exploration of the effects of Lorentz violation in both de Sitter and anti-de Sitter spacetimes. Given the accelerated expansion of the universe, current models suggest that the universe is asymptotically de Sitter. Thus, this paper aims to explore the effects of Lorentz violation on the dynamical properties of black holes in an asymptotically de Sitter spacetime. Furthermore, within this framework, more black hole solutions have been solved Filho et al. (2023); Övgün et al. (2019); Ding et al. (2020, 2023, 2022). Since then, this solution has served as a foundation for further research into related topics, including quasinormal modes (QNMs) Oliveira et al. (2021); Liu et al. (2023a); Malik (2024); Mai et al. (2024), black hole shadows Jha and Rahaman (2021); Kuang and Övgün (2022); Gao (2024a); Pantig et al. (2024); Liu et al. (2024c, d, e); Afrin et al. (2024), photon trajectories Lambiase et al. (2023a), and studies on entanglement degradation Liu et al. (2025a, b) or particle creation Araújo Filho (2025), as well as the metric affine case Filho et al. (2024b); Amarilo et al. (2024); Lambiase et al. (2023b); Araújo Filho et al. (2024b); Gao (2024b); Heidari et al. (2024); Jha and Rahaman (2024), among others Araújo Filho et al. (2025); Filho et al. (2024c); Liu et al. (2024f); Ji et al. (2024).
@@
On the other hand, in black hole physics, interactions between black holes and matter fields are crucial for understanding black hole stability, testing the no-hair theorem Bhattacharya and Lahiri (2007), and developing new black hole solutions Herdeiro et al. (2016). To study these interactions, researchers often use the linear approximation of black hole perturbation theory Calcagni and Modesto (2017); Mukohyama et al. (2022). Black hole perturbations can be divided into two types. The first type involves perturbations of the black hole’s metric itself, where gravitational radiation from black hole disturbances exceeds that caused by external fields Del Águila and Matos (2021), making it a key measure of black hole stability. The second type involves perturbations within the ambient spacetime of the black hole, arising from disturbances caused by matter fields. These perturbations reflect the dynamic behavior of matter fields near the black hole and offer insights into the interactions between black holes and their environment. Similar to ripples on water caused by external forces, spacetime oscillations propagate as gravitational waves, with energy gradually dissipating during transmission. Perturbative fields in black hole spacetimes typically undergo three stages: initial bursts, QNMs, and power-law tails Destounis et al. (2020). The latter two stages, linked to the intrinsic properties of the black hole, are of particular interest. These stages, often described as the ”ringdown” of black holes, exhibit characteristics akin to the acoustic signatures of black holes Chakrabarti and Tchrakian (2002). By analyzing these signatures, researchers can identify and categorize different types of black holes.
@@
This paper investigates the QNMs and the dynamic waveform evolution in de Sitter black hole spacetimes within the Einstein—Bumblebee gravitational model and contrasts these outcomes with the results from traditional Schwarzschild-de Sitter spacetimes Zhidenko (2004). The impact of different Lorentz violation parameters on the properties of black hole spacetimes is explored. Sec. II of this manuscript provides a brief overview of the Einstein—Bumblebee gravity model and its underlying geometrical background and derives equations of motion for the scalar field. Sec. III describes the WKB method Konoplya (2003), employs numerical techniques to find the scalar field QNMs, and presents the associated numerical results. Sec. III briefly introduces the matrix method Lin et al. (2017) and the WKB method Konoplya (2003) to solve the quasinormal modes of the scalar field and presents the corresponding numerical calculation results. In Sec. IV, we delve into the dynamical evolution of the scalar field under a small cosmological constant, revealing significant late-time behavior characterized by power-law tails. The final section offers a perspective on potential future research directions for this topic.
II Einstein—Bumblebee Gravity Model
@@
Here, $T_{\mu\nu}^{M}$ is the energy—momentum tensor of the matter field and
@@
is the Bumblebee energy—momentum tensor with
@@
By solving this system of differential equations, one can obtain curved spacetime metrics with asymptotically de Sitter (dS) or anti-de Sitter (AdS) properties under the influence of the Bumblebee field. These can also be referred to as Einstein—Bumblebee black hole metrics, and their line element is given by:
@@
The metric function in references can thus be rewritten as $F(r)=1-\frac{2M}{r}-\frac{\Lambda r^{2}}{3}$, which is consistent with the Schwarzschild—de Sitter case. However, the asymptotic behavior of the metric at infinity is not consistent with the Schwarzschild—de Sitter case, and the influence of Lorentz violation also needs to be considered. When the cosmological constant $\Lambda$ is zero, the resulting metric coincides with the results obtained by Casana in 2018 Casana et al. (2018). When the Lorentz violation parameter $\ell=0$, the metric simplifies to Schwarzschild—(A)dS. The incorporation of the cosmological constant modifies the structure of spacetime, resulting in a dual horizon configuration that encompasses both the black hole event horizons and cosmological horizons. This means that, the metric function can be reformulated as Jing (2004); Liu et al. (2023b, 2024g):
@@
II.1 Massive Scalar Field Perturbation Equation
The equation of motion for a massive scalar field is described by the Klein—Gordon equation:
$\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Phi\right)=\mu^{2}\Phi,$ (9)
@@ II.1 Massive Scalar Field Perturbation Equation
$V(r)=$ $\frac{F(r)}{r^{2}}\left[(1+\ell)L(L+1)+(1+\ell)r^{2}\mu^{2}+\frac{2M}{r}-\frac{2\Lambda r^{2}}{3}\right].$ (13)
@@
We can use standard methods to solve the eigenfrequencies of the equation, such as the matrix method Lin and Qian (2017); Lin et al. (2017); Lin and Qian (2019). It is worth noting that if the field mass is not considered, we can also use the WKB method to verify the results Iyer and Will (1987). Meanwhile, we plotted the potential $V(r)$ against $r$ for various parameters in Fig. 1. From the potential plot, when $\mu=0$, we can see that it satisfies the criteria required by the WKB method.
@@
In such black hole spacetime systems, the dissipation of energy primarily arises from the imposition of fully absorbing boundary conditions. Specifically, this includes two requirements: at the black hole horizon, the perturbation must be an incoming wave, and at infinity, it should be an outgoing wave. These boundary conditions apply to asymptotically flat spacetimes. More precisely, the boundary conditions encompass two stipulations: perturbations must manifest as ingoing waves at the black hole horizon and as outgoing waves at infinity. However, in de Sitter or anti-de Sitter spacetimes, the choice of boundary conditions differs. In de Sitter spacetimes, the de Sitter horizon replaces the role traditionally played by spatial infinity; in anti-de Sitter spacetimes, the requirement for the convergence of solutions at spatial infinity supersedes the condition for outgoing waves. These distinct boundary conditions reflect the unique properties of different types of spacetimes. If equations in the frequency domain are considered, one can separate variables in the wave function, allowing for $\psi(t,r)=\psi(r)e^{-i\omega t}$, the socalled harmonic ansatz, leading to
$\frac{d^{2}}{dr_{*}^{2}}\psi(r)+\left[(1+\ell)\omega^{2}-V(r)\right]\psi(r)=0.$ (17)
The quasinormal mode problem fundamentally represents a frequency eigenvalue problem for the wave equation. The real and imaginary parts of the quasinormal mode frequency $\omega=\omega_{R}+i\omega_{I}$ each have distinct physical meanings expressed by $e^{-i\omega t}=e^{-i\omega_{R}t}e^{\omega_{I}t}$. The real part $\omega_{R}$ clearly represents the oscillation frequency of the perturbation mode, whereas the imaginary part $\omega_{I}$ indicates the rate of exponential growth or decay of the perturbation. If $\omega_{I}$ is negative, it indicates that the perturbation decays over time, suggesting that the mode is stable; conversely, a positive $\omega_{I}$ signifies an unstable mode, under which the black hole spacetime might undergo substantial changes.
@@ III Numerical Methods and Quasinormal Modes
To determine the frequencies of quasinormal modes, one must impose physically meaningful boundary conditions on the solutions of the wave equation in black hole spacetimes. These conditions dictate the asymptotic behavior of the scalar field, which varies depending on the spacetime geometry. In general, the field’s asymptotic form is expressed as:
$\psi(r)\sim\begin{cases}e^{G(\omega)r_{*}},&r_{*}\to-\infty,\\ r^{\alpha}e^{K(\omega)r_{*}},&r_{*}\to\infty,\end{cases}$ (18)
where $G(\omega)$ and $K(\omega)$ are functions governed by the spacetime’s asymptotic properties, specifically the effective potential near the boundaries, and $\alpha$ reflects the physical system’s characteristics Dolan (2007); Corrêa et al. (2024). In Einstein—Bumblebee black holes with massive fields, for example, we set $G(\omega)=-i\sqrt{1+\ell}\omega$ and $K(\omega)=\pm\sqrt{(1+\ell)(\omega^{2}-\mu^{2})}$. Here, solutions with $\text{Re}[K(\omega)]>0$ represent quasinormal modes with outgoing behavior at infinity, while those with $\text{Re}[K(\omega)]<0$ indicate exponentially decaying quasibound states, the latter being outside this study’s scope Dolan (2007). These distinct spectra coexist, each influencing the field’s evolution differently, with quasibound states notably linked to superradiant instabilities. By contrast, in Einstein-Bumblebee de Sitter spacetime, where the effective potential diminishes at both the event and cosmological horizons, we adopt $G(\omega)=-i\sqrt{1+\ell}\omega$ and $K(\omega)=i\sqrt{1+\ell}\omega$, yielding oscillatory solutions at both limits:
$\psi(r)\sim\begin{cases}e^{-i\sqrt{1+\ell}\omega r_{*}},&r_{*}\to-\infty,\\ e^{i\sqrt{1+\ell}\omega r_{*}},&r_{*}\to\infty,\end{cases}$ (19)
irrespective of the field’s mass, thus eliminating the quasibound spectrum. In what follows, we will explore three distinct approaches to calculate the quasinormal frequencies in Einstein—Bumblebee de Sitter black holes.
III.1 Matrix Method
The Matrix Method (MM) offers a versatile approach to solving eigenvalue problems by employing an interpolation strategy that avoids reliance on a structured grid. This technique transforms the perturbation equations, together with their boundary constraints, into a homogeneous matrix equation through the discretization of the underlying linear partial differential equation Lin et al. (2017). At the core of this method, the unknown eigenfunction is approximated using Taylor series expansions, constructed up to the $N$th order around $N$ selected points within the domain. These points, while capable of being irregularly spaced to optimize precision versus computational cost, are in this instance chosen to form a uniform grid over $[0,1]$. Solving the resulting system of linear algebraic equations yields the desired eigenvalue.
To simplify the radial domain, a change of variable is introduced:
$x=\frac{r-r_{h}}{r_{c}-r_{h}},$ (20)
mapping the interval $r\in[r_{h},r_{c}]$, which represents the outer communication region of a black hole, onto $x\in[0,1]$. The function $\chi(x)$ is defined to vanish at the boundaries, i.e., $\chi(0)=0$ and $\chi(1)=0$, and is linked to the wave function $\psi_{l}^{\pm}(x)$ via:
$\psi_{l}^{\pm}(x)=x(1-x)(1-x)^{i\sqrt{(1+\ell)}\omega\eta_{c}}x^{-i\sqrt{1+\ell}\omega\eta_{h}}\chi(x).$ (21)
The perturbation equations are then reexpressed in a compact differential form:
$\tau_{0}^{\pm}(x)\chi^{\prime\prime}(x)+\lambda_{0}^{\pm}(x)\chi^{\prime}(x)+s_{0}^{\pm}(x)\chi(x)=0,$ (22)
where the coefficients $\tau_{0}^{\pm}(x)$, $\lambda_{0}^{\pm}(x)$, and $s_{0}^{\pm}(x)$ are numerically evaluated based on the effective potentials, the coupling parameter $\xi$, and the black hole horizons $r_{h}$ and $r_{c}$. These horizons depend on the mass $M$, the effective cosmological constant $\Lambda_{e}$, and the multipole number $\ell$.
For numerical solution, the differential equation is discretized across a uniform set of grid points in $[0,1]$. Around each point, $\chi(x)$ is expanded via Taylor series, enabling the construction of differential matrices. This process converts the differential equation into a matrix system:
$\left(M^{\pm}_{0}+\omega M_{1}^{\pm}\right)\chi(x)=0,$ (23)
where $\mathbf{\chi}$ is a vector containing the values of $\chi(x)$ at the grid points, and $M_{0}^{\pm}$ and $M_{1}^{\pm}$ are matrices built from the coefficients and differential operators. The QNM frequencies $\omega$ are determined by finding the values that satisfy:
$|M^{\pm}_{0}+\omega M_{1}^{\pm}|=0.$ (24)
A distinguishing feature of the MM is its adaptability: the placement of interpolation points can be tailored to enhance accuracy in critical regions, though here a uniform distribution is adopted for simplicity. This flexibility makes the method particularly effective for balancing computational efficiency with the precision required in complex physical systems like black hole perturbation analyses.
III.2 Wenzel—Kramers—Brillouin (WKB) Approximation Method